OpenStudy (anonymous):

Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

what

OpenStudy (anonymous):

@nirmalnema

OpenStudy (nirmalnema):

let x be 40% of 0.5 x=0.5*40 /100 x=0.2 they have to earn total profit of $500 the require to sell total no of flowers be y. they are getting a profit of 0.2 on each flower 500/0.2 be total flowers sold =2500

OpenStudy (nirmalnema):

do u get that..!! @yololifelove

OpenStudy (anonymous):

yes thank u :)